OpenStudy (anonymous):

Solve –3x2 + 30x – 90 = 0

OpenStudy (abb0t):

factor out the 3 -3(x^2-10x+30) Use quadratic formula to solve next.

OpenStudy (abb0t):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

The Answer I Got Was \[X=5\pm 2i \sqrt{5} \]

OpenStudy (anonymous):

@abb0t

hero (hero):

–3x2 + 30x – 90 = 0 3x^2 - 30x + 90 = 0 3(x^2 - 10x + 30) = 0 x^2 - 10x + 30 = 0 That should make it easier

OpenStudy (anonymous):

@hero so is my answer correct

OpenStudy (abb0t):

Looks about right, haven't solved for it. You can leave the "i" inside the root.